OpenStudy (anonymous):

Help Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (4 points) Part B: Write the equation obtained in Part A using function notation. (2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

Part A: Well, each day it costs $93 to rent. So, y= x*93.

OpenStudy (radar):

Note that does (x*93) not work for the 7 day rental of $625. Need to do some more thinking.

OpenStudy (radar):

What are you now studying, "inequalities" ?

OpenStudy (radar):

I don't know what you are studying, maybe step function ???

OpenStudy (radar):

This is just for comment, a feeble attempt to capture this problem. a swag lol

OpenStudy (radar):

\[y =(0< x <)\times $93 +(\le x)\times $89.30\]

OpenStudy (radar):

\[y=(0<x <7)*$93 + (7\le x)\times $89.30\]

OpenStudy (radar):

Omitted 7 in the first one. Ignore the first try.